Find the inverse function using diagrams and test points.
1: f(x) = 2x + 1
2: f(x) = 4x – 2
3: f(x) = sqrt(x-8)

Answer by rapture(86)

Tip: DO NOT POST MORE THAN ONE QUESTION PER POST OR YOUR PROBLEMS WILL BE IGNORED.


I will do question 1 only.


f(x) = 2x + 1


Let f(x) = y


y = 2x + 1


Switch y with x.


x = 2y + 1


Solve for y.


x - 1 = 2y


(x - 1)/2 = y


Replace y with the inverse function notation.


(x - 1)/2 = f^(-1) x


Do the rest in similar fashion.


harpazo1965@gmail.com